# Linear Quadratic Control with Non-Markovian and Non-Semimartingale Noise Models

### v1

@@ I Introduction
Linear quadratic Gaussian (LQG) optimal control is a fundamental problem in stochastic control theory and widely used in various engineering applications. The standard LQG framework assumes that the system dynamics and observations are corrupted by Brownian noise processes allowing for the derivation of a separation principle. The optimal controller can then be obtained through the application of Kalman filtering for optimal state estimation and using the Hamilton-Jacobi-Bellman (HJB) equation to derive the optimal feedback control. The analytical machinery for each of these relies on the standard Itô stochastic calculus. For theoretical background on the standard LQG problem, we refer the reader to [book4].
@@
However, empirical measurements of various physical phenomena often indicate that Brownian processes are not always the most effective modeling choice. Many real-world processes exhibit dependencies, correlations, and sample paths that have low Hölder regularity, which may not be adequately modeled by standard Brownian processes, and instead require non-Markovian and/or non-semimartingale models. For instance, due to its ability to capture memory effects and self-similarity properties, empirically, fractional Brownian motion (fBm) has been shown to better model a broad range of phenomena such as in financial systems [ted], turbulent systems [turbulence1, turbulence2], rainfall systems [rainfall], and biological systems [biology]. Other examples include financial data modeled as stable Lévy processes [financial] and noise processes in quantum systems modeled as shot noise [quantum].
@@
Moving from Brownian processes to these more general processes in an estimation and control context poses theoretical challenges. Unlike Brownian motion, more general processes that may be non-Gaussian, non-Markovian and/or non-semimartingale require more analysis, such as careful characterization of more attributes like tail distributions, skewness, higher order moments, and dependence and correlation parameters. Ignoring these effects and forcing LQG controllers that rely on the noise processes being modeled as Brownian processes can significantly impact system behavior and control performance, even leading to system failures [qz].
@@
The extension of LQ control to systems affected by noise sequences that are more general than Brownian noise processes is, thus, an important area of research. One model that has seen significant work is when the noise is still Gaussian; however, instead of white Brownian noise, we consider a colored noise sequence. In this context, [bet] investigates stochastic LQR optimal control with both white and colored Gaussian noises, employing dynamic programming techniques to derive optimal solutions. A traditional approach here is to transform the colored noise into white noise through appropriate filtering, thereby allowing the application of standard LQG methods. Some recent studies considering such models include [jml], which proposes the use of shaping filters for systems with colored measurement noises, [jh], which explores how pink noise, a type of colored noise, affects optimal policies in reinforcement learning (highlighting the deviations from assumption of a white noise), and [fas], which demonstrates the application of optimal vibration control in the presence of colored noise, showcasing practical implications of such models.
@@
There is a line of work that studies control design with non-Markovian, but semimartingale noise, which allows the use of standard Itô stochastic calculus. For example, [as] addressed detection and estimation problems in heavy-tailed symmetric alpha-stable distributed noise, [ted3] addressed stochastic control in the presence of Rosenblatt process noise, [st] provided the necessary conditions for optimal control of stochastic systems with Markovian random jump noise models, and [yh] extended this work to partial information LQ control for jump diffusions, which can have non-Markovian coefficients for drift, diffusion, and jump terms.
@@
A particular noise process that has attracted much attention is fractional Brownian motion (fBm), which is a Gaussian process that is, in general, non-Markovian, as seen in the works of [mlk][tl] and [mlk2]. When the Hurst index exceeds 0.5, classical Itô stochastic calculus needs to be extended to the use of fractional stochastic calculus in these papers. Fractional calculus deals with derivatives and integrals of arbitrary, non-integer orders [bg]. Several works have studied optimal LQR control design for a system driven by fBm, including [ted][ted2][yh2] and [ted4].
@@
When considering low Hölder regularity non-semimartingale noise sequences, the lack of an Itô integral necessitates alternative mathematical frameworks to be able to analyze systems driven by such noise sequences. To analyze such systems, we employ rough path theory, which encompasses and generalizes classical stochastic calculus. Rough path theory provides a toolbox for the analysis of stochastic differential equations driven by non-semimartingale processes and rough signals that have low Hölder regularity [book, book2, book3][jd] provided a foundational contribution by exploring stochastic control using rough path theory when considering standard Brownian motion noise (i.e. Hölder regularity $H=1/2$). That paper formulated controlled rough stochastic differential equations and demonstrated that the value function satisfies a Hamilton-Jacobi-Bellman (HJB) type equation, and provides a continuous map for the solution. This continuity property may not, in general, hold with standard Itô calculus. Thus, rough path analysis can provide a continuous pathwise deterministic solution to stochastic differential equations driven by rough signals. In addition, in the case of fractional noises such as fBm, fractional Stratonovich integrals can be approximated using Wong-Zakai and mollifier approximations, as discussed by [ct], for rough differential equations.
@@ II Rough Path Theory Preliminaries
Definition 1 (Rough Paths [book]).
@@
Definition 2 (Hölder Regularity of a Rough Path [book]).
@@
is well defined under the condition that the path $x^{\prime}(t)$ is unique. The uniqueness holds if one of the following conditions are satisfied [gubinelli]:
@@
admits a unique solution [lyons]. Due to the use of Lyons’ rough path theory framework, we consider classes of stochastic processes that have rough path lifts, and the lifts are well defined using Lyons’ rough path framework. Those include classical stochastic processes such as semimartingales and Markov processes, as well as other stochastic processes such as fBm [lyons].
@@ IV-A Optimal Controller with State Observation
Lemma 1 provides the definition of the solution of the rough HJB equation using the classical HJB equation [allan].
Lemma 1 (Definition 3.13 in [allan]).
@@
Using the variation of constants formula [book, Chapter 5.3], we can write
@@
is continuous and the integrand is decaying due to Assumption 1. Thus, there exist constants $C>0$ and $\alpha\in(0,1]$ such that, as shown in [book3],
@@
Corollary 3 formally proves the pathwise optimality of the optimal controller in Theorem 2. This is a generalization of the pathwise optimality in [jd], where standard Brownian motion is considered.
@@
Under Assumption 1, the algebraic Riccati equation has a unique symmetric positive definite solution P. Thus, the fundamental solution $\boldsymbol{\Phi}(s,t)$ for (7) exists. Hence, the rough integral $V_{noise}$ in (8) is well defined for rough path v [gubinelli].
@@ IV-B Optimal Controller without State Observation
under the assumption that $v(t)$ and $w(t)$ are uncorrelated with $e(t)$ for optimal estimation, where $\Sigma_{v}=\mathbb{E}[d\mathbf{v}(t)d\mathbf{v}(t)^{\top}]$, $\Sigma_{w}=\mathbb{E}[d\mathbf{w}(t)d\mathbf{w}(t)^{\top}]$, $\textbf{R}_{vw}=\mathbb{E}[d\mathbf{v}(t)d\mathbf{w}(t)^{\top}]$ and $\textbf{R}_{wv}=\mathbb{E}[d\mathbf{w}(t)d\mathbf{v}(t)^{\top}]$ that are covariance and cross-correlation operators, respectively. They are interpreted as covariance and cross-correlation kernels, or integral operators [book3, operators]. The interchange between the time derivative and expectation operators is a result of the dominated convergence theorem and the rough Itô formula [book]. Also, the second order terms in time, i.e. $(dt)^{2}$, in $\mathbb{E}\big[de(t)de(t)^{\top}\big]$ can be ignored since they are equal to 0.
@@
Furthermore, it is known that, in general, Itô solution maps are not continuous [book2]. On the contrary, the rough path framework provides solution maps that are locally uniformly continuous [book2]. This provides another motivation for the use of rough path theory over classical stochastic calculus for the case when both solutions are well defined. Theorem 5 formally states this continuity result.
@@
    are, in general, not continuous in supremum norm nor in any other metric on path space, since they depend on stochastic integrals whose values can jump under noise perturbations [book2, lyons, lyons2, friz].
@@
The proof of the discontinuity of the Itô solution follows from Theorem 1.1.1 in [book2].
@@
By [continuity], the filter solution mapping that associates the solution of the stochastic differential equation to an enhanced observation path is locally uniformly continuous. Concretely, Theorem 2 (existence of solution) and Theorem 3 (continuity) in [continuity] provide the detailed proof for the continuity of the mapping, which is the rough path replacement of the Clark robust representation in [clark]. This means that the solution of
@@ Appendix A Supporting Theorems and Lemmas
Lemma 6 (Sewing Lemma in [book]).
@@
Lemma 7 (Lyons’ Universal Limit Theorem in [book2]).

### v2

@@ I Introduction
Linear quadratic Gaussian (LQG) optimal control is a fundamental problem in stochastic control theory and widely used in various engineering applications. The standard LQG framework assumes that the system dynamics and observations are corrupted by Brownian noise processes allowing for the derivation of a separation principle. The optimal controller can then be obtained through the application of Kalman filtering for optimal state estimation and using the Hamilton-Jacobi-Bellman (HJB) equation to derive the optimal feedback control. The analytical machinery for each of these relies on the standard Itô stochastic calculus. For theoretical background on the standard LQG problem, we refer the reader to [2].
@@
However, empirical measurements of various physical phenomena often indicate that Brownian processes are not always the most effective modeling choice. Many real-world processes exhibit dependencies, correlations, and sample paths that have low Hölder regularity, which may not be adequately modeled by standard Brownian processes, and instead require non-Markovian and/or non-semimartingale models. For instance, due to its ability to capture memory effects and self-similarity properties, empirically, fractional Brownian motion (fBm) has been shown to better model a broad range of phenomena such as in financial systems [12], turbulent systems [27, 28], rainfall systems [24], and biological systems [36]. Other examples include financial data modeled as stable Lévy processes [35] and noise processes in quantum systems modeled as shot noise [7].
@@
Moving from Brownian processes to these more general processes in an estimation and control context poses theoretical challenges. Unlike Brownian motion, more general processes that may be non-Gaussian, non-Markovian and/or non-semimartingale require more analysis, such as careful characterization of more attributes like tail distributions, skewness, higher order moments, and dependence and correlation parameters. Ignoring these effects and forcing LQG controllers that rely on the noise processes being modeled as Brownian processes can significantly impact system behavior and control performance, even leading to system failures [41].
@@
The extension of LQ control to systems affected by noise sequences that are more general than Brownian noise processes is, thus, an important area of research. One model that has seen significant work is when the noise is still Gaussian; however, instead of white Brownian noise, we consider a colored noise sequence. In this context, [16] investigates stochastic LQR optimal control with both white and colored Gaussian noises, employing dynamic programming techniques to derive optimal solutions. A traditional approach here is to transform the colored noise into white noise through appropriate filtering, thereby allowing the application of standard LQG methods. Some recent studies considering such models include [31], which proposes the use of shaping filters for systems with colored measurement noises, [21], which explores how pink noise, a type of colored noise, affects optimal policies in reinforcement learning (highlighting the deviations from assumption of a white noise), and [37], which demonstrates the application of optimal vibration control in the presence of colored noise, showcasing practical implications of such models.
@@
There is a line of work that studies control design with non-Markovian, but semimartingale noise, which allows the use of standard Itô stochastic calculus. For example, [38] addressed detection and estimation problems in heavy-tailed symmetric alpha-stable distributed noise, [15] addressed stochastic control in the presence of Rosenblatt process noise, [39] provided the necessary conditions for optimal control of stochastic systems with Markovian random jump noise models, and [22] extended this work to partial information LQ control for jump diffusions, which can have non-Markovian coefficients for drift, diffusion, and jump terms.
@@
A particular noise process that has attracted much attention is fractional Brownian motion (fBm), which is a Gaussian process that is, in general, non-Markovian, as seen in the works of [25][29] and [26]. When the Hurst index exceeds 0.5, classical Itô stochastic calculus needs to be extended to the use of fractional stochastic calculus in these papers. Fractional calculus deals with derivatives and integrals of arbitrary, non-integer orders [19]. Several works have studied optimal LQR control design for a system driven by fBm, including [12][13][23] and [14].
@@
When considering low Hölder regularity non-semimartingale noise sequences, the lack of an Itô integral necessitates alternative mathematical frameworks to be able to analyze systems driven by such noise sequences. To analyze such systems, we employ rough path theory, which encompasses and generalizes classical stochastic calculus. Rough path theory provides a toolbox for the analysis of stochastic differential equations driven by non-semimartingale processes and rough signals that have low Hölder regularity [17, 33, 18][11] provided a foundational contribution by exploring stochastic control using rough path theory when considering standard Brownian motion noise (i.e. Hölder regularity $H=1/2$). That paper formulated controlled rough stochastic differential equations and demonstrated that the value function satisfies a Hamilton-Jacobi-Bellman (HJB) type equation, and provides a continuous map for the solution. This continuity property may not, in general, hold with standard Itô calculus. This has been followed up in [30] where a rough stochastic Pontryagin principle was derived, as well as optimal relaxed control theory with rough differential equations in [4, 8]. In all of these settings, rough path analysis can provide a continuous pathwise deterministic solution to stochastic differential equations driven by rough signals. In addition, in the case of fractional noises such as fBm, fractional Stratonovich integrals can be approximated using Wong-Zakai and mollifier approximations, as discussed by [40], for rough differential equations. More recently there has also been significant interest in exploiting signature-based methods for optimal control [6, 5]. Nonetheless, very little attention has been paid to control design questions.
@@ II Rough Path Theory Preliminaries
Definition 1 (Rough Paths [17]).
@@
Definition 2 (Hölder Regularity of a Rough Path [17]).
@@
is well defined under the condition that the path $x^{\prime}(t)$ is unique. The uniqueness holds if one of the following conditions are satisfied [20]:
@@
admits a unique solution [32]. Due to the use of Lyons’ rough path theory framework, we consider classes of stochastic processes that have rough path lifts, and the lifts are well defined using Lyons’ rough path framework. Those include classical stochastic processes such as semimartingales and Markov processes, as well as other stochastic processes such as fBm [32].
@@ IV-A Optimal Controller with State Observation
Lemma 1 provides the definition of the solution of the rough HJB equation using the classical HJB equation [1].
Lemma 1 (Definition 3.13 in [1]).
@@
Using the variation of constants formula [17, Chapter 5.3], we can write
@@
is continuous and the integrand is decaying due to Assumption 1. Thus, there exist constants $C>0$ and $\alpha\in(0,1]$ such that, as shown in [18],
@@
Corollary 3 formally proves the pathwise optimality of the optimal controller in Theorem 2. This is a generalization of the pathwise optimality in [11], where standard Brownian motion is considered.
@@
Under Assumption 1, the algebraic Riccati equation has a unique symmetric positive definite solution P. Thus, the fundamental solution $\boldsymbol{\Phi}(s,t)$ for (7) exists. Hence, the rough integral $V_{noise}$ in (8) is well defined for rough path v [20].
@@ IV-B Optimal Controller without State Observation
under the assumption that $v(t)$ and $w(t)$ are uncorrelated with $e(t)$ for optimal estimation, where $\Sigma_{v}=\mathbb{E}[d\mathbf{v}(t)d\mathbf{v}(t)^{\top}]$, $\Sigma_{w}=\mathbb{E}[d\mathbf{w}(t)d\mathbf{w}(t)^{\top}]$, $\textbf{R}_{vw}=\mathbb{E}[d\mathbf{v}(t)d\mathbf{w}(t)^{\top}]$ and $\textbf{R}_{wv}=\mathbb{E}[d\mathbf{w}(t)d\mathbf{v}(t)^{\top}]$ that are covariance and cross-correlation operators, respectively. They are interpreted as covariance and cross-correlation kernels, or integral operators [18, 3]. The interchange between the time derivative and expectation operators is a result of the dominated convergence theorem and the rough Itô formula [17]. Also, the second order terms in time, i.e. $(dt)^{2}$, in $\mathbb{E}\big[de(t)de(t)^{\top}\big]$ can be ignored since they are equal to 0.
@@
Furthermore, it is known that, in general, Itô solution maps are not continuous [33]. On the contrary, the rough path framework provides solution maps that are locally uniformly continuous [33]. This provides another motivation for the use of rough path theory over classical stochastic calculus for the case when both solutions are well defined. Theorem 5 formally states this continuity result.
@@
    are, in general, not continuous in supremum norm nor in any other metric on path space, since they depend on stochastic integrals whose values can jump under noise perturbations [33, 32, 34, 18].
@@
The proof of the discontinuity of the Itô solution follows from Theorem 1.1.1 in [33].
@@
By [10], the filter solution mapping that associates the solution of the stochastic differential equation to an enhanced observation path is locally uniformly continuous. Concretely, Theorem 2 (existence of solution) and Theorem 3 (continuity) in [10] provide the detailed proof for the continuity of the mapping, which is the rough path replacement of the Clark robust representation in [9]. This means that the solution of
@@ Appendix A Supporting Theorems and Lemmas
Lemma 6 (Sewing Lemma in [17]).
@@
Lemma 7 (Lyons’ Universal Limit Theorem in [33]).